# Foundation Models in Smart Agriculture: Basics, Opportunities, and Challenges

### 21 Dec 2023 v3

@@ Abstract
The past decade has witnessed the rapid development and adoption of machine learning (ML) and deep learning (DL) methodologies in agricultural systems, showcased by great successes in applications such as weed control, plant disease detection, agricultural robotics, and precise livestock management. However, these conventional ML/DL models have certain limitations: They heavily rely on large, costly-to-acquire labeled datasets for training, require specialized expertise for development and maintenance, and are mostly tailored for specific tasks, thus lacking generalizability. Recently, large pre-trained models, also known as foundation models (FMs), have demonstrated remarkable successes in language, vision, and decision-making tasks across various domains. These models are trained on a vast amount of data from multiple domains and modalities. Once trained, they can accomplish versatile tasks with just minor fine-tuning and minimal task-specific labeled data. Despite their proven effectiveness and huge potential, there has been little exploration of applying FMs to agriculture artificial intelligence (AI). Therefore, this study aims to explore the potential of FMs in the field of smart agriculture. In particular, we present conceptual tools and technical background to facilitate the understanding of the problem space and uncover new research directions in this field. To this end, we first review recent FMs in the general computer science (CS) domain and categorize them into four categories: language FMs, vision FMs, multimodal FMs, and reinforcement learning FMs. Subsequently, we outline the process of developing agriculture FMs (AFMs) and discuss their potential applications in smart agriculture. We also discuss the unique challenges and risks associated with developing AFMs, including model training, validation, and deployment. Through this study, we contribute to the advancement of AI in agriculture by introducing AFMs as a promising paradigm that can significantly mitigate the reliance on extensive labeled datasets and enhance the efficiency, effectiveness, and generalization of agricultural AI systems. To facilitate further research, a well-classified and actively updated list of papers on AFMs are organized and accessible at https://github.com/JiajiaLi04/Agriculture-Foundation-Models.
@@ 1 Introduction
Figure 1: Overview of building foundation models for agricultural applications.
@@
Modern agricultural systems have been undergoing a revolution through the tight integration of recent advances in information and communication technologies (ICT). This has led to a prosperous field known as smart farming (also referred to as smart agriculture) (Gondchawar et al., 2016, Wolfert et al., 2017, Walter et al., 2017), showing promise in enhancing agricultural outputs, increasing farming efficiency, and improving the quality of the final product. These technologies feature multi-disciplinary advancements involving unmanned aerial/ground vehicles (UAVs/UGVs), image processing, machine learning, big data, cloud computing, and wireless sensor networks (WSNs) (Walter et al., 2017, Wolfert et al., 2017, Moysiadis et al., 2021), enabling farmers to make well-informed decisions regarding planting, tending, and harvesting to maximize productivity and profits. However, efficient extraction of relevant information from diverse data sources, especially imaging data, presents a significant challenge. Traditional data mining techniques often struggle to uncover meaningful insights from such complex data (Wolfert et al., 2017).
@@
Deep learning (DL) (LeCun et al., 2015), on the other hand, has demonstrated remarkable capabilities in processing complex, high-dimensional data and has achieved numerous successes in a myriad of applications. DL methods excel in feature extraction, pattern classification, and learning high-quality image representations, showcasing promising performance across various agricultural domains (Kamilaris and Prenafeta-Boldú, 2018). These applications include weed control (Hasan et al., 2021, Chen et al., 2022), plant disease detection (Mohanty et al., 2016, Xu et al., 2021), plant phenotyping (Li et al., 2020, Xiang and Wang, 2023), postharvest quality assessment (Mendigoria et al., 2021, Zhou et al., 2022), and robotic fruit harvesting (Chu et al., 2021, Zhang et al., 2022a, Chu et al., 2023), among others. Despite the progress, these approaches heavily rely on supervised training, which hinges on large-scale, task-specific, and high-quality labeled datasets (Sun et al., 2017). However, collecting and annotating such datasets is extremely time-consuming, resource-intensive, and expensive, creating a high bar for resource-limited applications (Lu and Young, 2020, Xu et al., 2023). This challenge is particularly pronounced in specific applications such as plant disease detection, weed recognition, and fruit defect detection, where constraints related to biological materials, imaging conditions, and the need for precise pixel-level annotations arise (Lu and Young, 2020). Moreover, the collected datasets are often limited in their generalizability to even similar agricultural domains or applications, necessitating repetition in the data collection and model development process (Ghazi et al., 2017, Thenmozhi and Reddy, 2019). This repetitive process not only adds to the overall time and monetary costs but also hinders the generalizability, efficiency, and scalability of DL frameworks in agricultural applications.
@@
To mitigate the aforementioned issues, many approaches have been proposed, including transfer learning (Weiss et al., 2016, Zhuang et al., 2020), few-shot learning (Wang et al., 2020, Yang et al., 2022a), label-efficient learning (Zhou, 2018, Li et al., 2023a), among others. Specifically, transfer learning has emerged as a widely adopted approach in agricultural applications and beyond (Espejo-Garcia et al., 2020, Chen et al., 2022, Abbas et al., 2021, Dang et al., 2022), which leverages DL models pre-trained on large-scale image datasets, such as ImageNet (Deng et al., 2009), Microsoft COCO (Lin et al., 2014), and PlantCLEF2022 (Goëau et al., 2022). The pre-trained models are then fine-tuned for specific tasks. For instance, Chen et al. (2022) evaluated 35 state-of-the-art DL models for weed classification using transfer learning. Few-shot learning, on the other hand, focuses on training models to quickly adopt new tasks or domains using only a limited number of labeled examples, by leveraging prior knowledge or meta-learning approaches. In contrast, label-efficient learning (Li et al., 2023a) employs weak supervision or no supervision techniques to mitigate the effects of laborious and time-consuming labeling challenges. However, these methods are often pre-trained on data from a single domain and modality, which restricts their adaptability to other domains and applications (Raffel et al., 2020). The above observations highlight the importance of incorporating diverse data sources to enhance their generalization capabilities across different domains and applications.
@@
Large pre-trained models also referred to as foundation models (FMs) (Bommasani et al., 2021, Moor et al., 2023, Mai et al., 2023), AI models designed to produce a wide and general variety of outputs, undergo extensive training on vast and diverse datasets, enabling them to tackle numerous downstream tasks, such as speech recognition and natural language understanding (Radford et al., 2019, Brown et al., 2020, OpenAI, 2022, Ouyang et al., 2022, Touvron et al., 2023), computer vision (Rombach et al., 2022, Ramesh et al., 2022, Kirillov et al., 2023, Kang et al., 2023), and decision making (Reed et al., 2022, Team et al., 2023). FMs are generally trained on large-scale datasets from diverse domains and modalities through self-supervised learning (Liu et al., 2021), and are capable of handling a wide range of applications from various domains with minimum fine-tuning and no/little task-specific labeled data (Bommasani et al., 2021, Moor et al., 2023, Mai et al., 2023). For example, the Generative Pre-trained Transformer 3 language model (GPT-3, (Brown et al., 2020)) demonstrated a breakthrough in in-context learning, enabling the model to perform novel tasks without explicit training, by leveraging text explanations (known as “prompts”) containing just a few examples. Subsequently, ChatGPT (OpenAI, 2022), a derivative of GPT-3, was developed and publicized by OpenAI for chat-based interactions and has brought about a transformative evolution in natural language processing, enabling unprecedentedly engaging conversational experiences. In the computer vision domain, FMs such as the segment anything model (SAM, (Kirillov et al., 2023)), trained on over 1 billion masks on 11M licensed and privacy-respecting images, can solve a range of downstream segmentation problems with new image distributions and tasks with zero-shot generalization. Deepmind Ada (Team et al., 2023), equipped with a customized Transformer architecture (Vaswani et al., 2017) using distillation with a teacher-student approach (Schmitt et al., 2018), brought FMs to reinforcement learning (RL) that generalizes to new tasks in a few-shot setting. Despite the aforementioned advances, FMs’ adoption in agriculture artificial intelligence (AI) has received scarce attention.
As such, this study aims to explore the potential of developing and applying FMs for agricultural applications. More specifically, we first review recent FMs (Section 2) in the general computer science (CS) domain and categorize them into four categories: language FMs, vision FMs, multimodal FMs, and reinforcement learning FMs. Subsequently, we outline the process of developing AFMs (Figure 1) and discuss their potential applications (Section 3) in smart agriculture. In addition, this study also identifies and discusses the unique challenges and potential pitfalls associated with developing AFMs (Section 4), including model training, validation, and deployment. This comprehensive examination of AFMs can be served as a valuable resource for newcomers to the field, as well as experienced researchers seeking to innovate within the agricultural space.
@@ 2 Foundation Models (FMs)
In this section, we first cover both the strengths and weaknesses of FMs, giving readers a balanced view of the topic. Then, we explore the different kinds of FMs, their distinct characteristics, and their applicability to various domains or tasks.
@@ 2.1 Pros and Cons of FMs
FMs represent a significant paradigm shift in the field of artificial intelligence (AI). In contrast to traditional AI models that are trained on domain-specific data for specialized tasks, FMs are pre-trained on a vast scale of data, spanning multiple domains and modalities (Bommasani et al., 2021, Moor et al., 2023, Mai et al., 2023). The training process allows these models to learn a broad base of knowledge and skills, much like the foundation of a building that supports all the structures above it. Once trained, these FMs can accomplish versatile tasks with just a little fine-tuning. This is analogous to the way humans transfer knowledge and skills from one area to another (Moor et al., 2023). They have led to breakthroughs in a variety of fields, from natural language processing to computer vision and beyond. We can summarize the pros of FMs as follows:
@@
Given the distinct advantages of FMs over traditional AI models, they exhibit substantial promise for agricultural applications. However, despite their progress, FMs are not without their unique set of challenges:
@@
For an in-depth discussion on deploying FMs in agricultural contexts, please refer to Section  4.
@@ 2.2 Categories of FMs
In this subsection, we present comprehensive reviews of FMs within the computer science (CS) domain, which were categorized into four distinct types (Figure 2) based on the primary data modality and learning paradigm they address: language FMs (LFMs), vision FMs (VFMs), multimodal FMs (MFMs), and reinforcement learning (RLFMs).
Figure 2: Timeline of foundation models (FMs).
@@ 2.2.1 Language foundation models (LFMs)
Over the past few years, the field of computational natural language has seen a complete transformation (Devlin et al., 2018, Brown et al., 2020) due to the rise of large-scale language modeling (LLMs). Essentially, language modeling (Nadkarni et al., 2011) involves predicting the subsequent element in a given sequence of tokens, relying on a self-supervised objective that requires no manual labeling beyond a raw text corpus.
The Bidirectional Encoder Representations from Trans-
formers (BERT, (Devlin et al., 2018)) is one of the most prominent models, with two versions: $\text{BERT}_{\text{BASE}}$ and
@@
$\text{BERT}_{\text{LARGE}}$, containing 110M and 340M parameters respectively. By simply incorporating an additional output layer for fine-tuning, the pre-trained BERT model can be adapted to generate cutting-edge models for a variety of tasks, such as question answering and language inference. Remarkably, this can be achieved without requiring substantial modifications to the task-specific architecture, even surpassing the performance of many dedicated task-specific architectures. Subsequently, Google proposed the Pathways Language Model (PaLM, (Wei et al., 2022)) and LLAMA (Touvron et al., 2023), further advancing the field of natural language processing. PaLM, with 540B parameters, was designed with a dense decoder-only Transformer model trained with the Pathways system (Barham et al., 2022) to enable efficient training. PaLM surpassed few-shot performance of prior large models, such as GLaM (Du et al., 2022), GPT-3 (Brown et al., 2020), and LLAMA (Touvron et al., 2023), on numerous very difficult tasks, such as nature language inference (Bowman et al., 2015), in-context reading comprehension (Jenkins et al., 2003), and question answering (Kwiatkowski et al., 2019).
@@
Participating in the development of large language models, OpenAI has proposed noteworthy contributions such as Generative Pretrained Transformer 2 (GPT-2, (Radford et al., 2019)), 3 (GPT-3, (Brown et al., 2020)), and 4 (GPT-4, (OpenAI, 2023)), having grown progressively more powerful, with 1.5 billion, 175 billion, and a staggering number of parameters, respectively, and setting remarkable benchmarks in language processing. ChatGPT (OpenAI, 2022), a specialized version of these models, has been extensively fine-tuned for human-like conversation. By adjusting the final output layer, the pre-trained GPT models can be tailored to deliver exceptional performance across a wide range of tasks, such as text generation, translation, and dialogue systems, outperforming many task-specific models without requiring extensive modifications to their foundational architecture. Recently, OpenAI showed that InstructGPT (Ouyang et al., 2022) with only 1.3B parameters are much better at following users’ instructions than GPT-3, by using an existing technique called reinforcement learning from human feedback (RLHF, (Christiano et al., 2017)).

### 17 Mar 2024 v4

# Large Language Models and Foundation Models in Smart Agriculture: Basics, Opportunities, and Challenges

@@ Abstract
The past decade has witnessed the rapid development and adoption of machine and deep learning (ML & DL) methodologies in agricultural systems, showcased by great successes in applications such as smart crop management, smart plant breeding, smart livestock farming, precision aquaculture farming, and agricultural robotics. However, these conventional ML/DL models have certain limitations: they heavily rely on large, costly-to-acquire labeled datasets for training, require specialized expertise for development and maintenance, and are mostly tailored for specific tasks, thus lacking generalizability. Recently, large pre-trained models, also known as foundation models (FMs), have demonstrated remarkable successes in language, vision, and decision-making tasks across various domains. These models are trained on a vast amount of data from multiple domains and modalities. Once trained, they can accomplish versatile tasks with just minor fine-tuning and minimal task-specific labeled data. Despite their proven effectiveness and huge potential, there has been little exploration of applying FMs to agriculture artificial intelligence (AI). Therefore, this study aims to explore the potential of FMs in the field of smart agriculture. In particular, conceptual tools and technical background are presented to facilitate the understanding of the problem space and uncover new research directions in this field. To this end, recent FMs in the general computer science (CS) domain are reviewed, and the models are categorized into four categories: language FMs, vision FMs, multimodal FMs, and reinforcement learning FMs. Subsequently, the process of developing agriculture FMs (AFMs) is outlined and their potential applications in smart agriculture are discussed. In addition, the unique challenges and risks associated with developing AFMs are discussed, including model training, validation, and deployment. Through this study, the advancement of AI in agriculture is explored by introducing AFMs as a promising paradigm that can significantly mitigate the reliance on extensive labeled datasets and enhance the efficiency, effectiveness, and generalization of agricultural AI systems. To facilitate further research, a well-classified and actively updated list of papers on AFMs is organized and accessible at https://github.com/JiajiaLi04/Agriculture-Foundation-Models.
@@ 1 Introduction
Figure 1: An overview of building FMs for agricultural applications. The pipeline depicts the process of (multi-modal) data collection and cleaning, training or fine-tuning of LLMs and FMs, and downstream adaptation. The terms “BERT” and “GPT-4” refer to Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2018) and Generative Pre-trained Transformer 4 (GPT-4) (Achiam et al., 2023), respectively.
@@
Modern agricultural systems have been undergoing a revolution through the tight integration of recent advances in information and communication technologies (ICT). This has led to a prosperous field known as smart farming (also referred to as smart agriculture) (Gondchawar et al., 2016, Wolfert et al., 2017, Walter et al., 2017), showing promise in enhancing agricultural outputs, increasing farming efficiency, and improving the quality of the final product. These technologies feature multi-disciplinary advancements involving unmanned aerial/ground vehicles (UAVs/UGVs), image processing, machine learning, big data, cloud computing, and wireless sensor networks (WSNs) (Walter et al., 2017, Wolfert et al., 2017, Moysiadis et al., 2021), enabling farmers to make well-informed decisions regarding planting, tending, and harvesting to maximize productivity and profits. However, efficient extraction of relevant information from diverse data sources, especially imaging data, presents a significant challenge. Traditional data mining approaches often struggle to uncover meaningful insights from such complex data (Wolfert et al., 2017). For instance, in agriculture, these techniques may face difficulties in extracting relevant features from datasets comprising diverse variables such as soil composition, weather patterns, crop health metrics, and geographical information (Wolfert et al., 2017).
@@
Deep learning (DL), on the other hand, has demonstrated remarkable capabilities in processing complex and high-dimensional data due to its ability to learn hierarchical representations of data (LeCun et al., 2015). Unlike traditional data mining techniques, which often rely on handcrafted feature extraction methods, DL models can automatically learn relevant features from raw data through multiple layers of abstraction. DL methods excel in feature extraction, pattern classification, and learning high-quality image representations, showing promising performance across various agricultural domains (Kamilaris and Prenafeta-Boldú, 2018). These applications include weed control (Hasan et al., 2021, Chen et al., 2022, Li et al., 2024, Chen et al., 2024), plant disease detection (Mohanty et al., 2016, Xu et al., 2021), plant counting (Pathak et al., 2022, Li et al., 2023c), plant phenotyping (Li et al., 2020, Xiang and Wang, 2023), postharvest quality assessment (Mendigoria et al., 2021, Zhou et al., 2022), robotic fruit harvesting (Chu et al., 2021, Zhang et al., 2022a, Chu et al., 2023), among others. Despite the progress, these approaches heavily rely on supervised training, which hinges on large-scale, task-specific, and high-quality labeled datasets (Sun et al., 2017). However, collecting and annotating such datasets is extremely time-consuming, resource-intensive, and expensive, thereby, creating a high bar for resource-limited applications (Lu and Young, 2020, Xu et al., 2023a). This challenge is particularly pronounced in specific applications such as plant disease detection, weed recognition, and fruit defect detection, where constraints related to biological materials, imaging conditions, and the need for precise annotations (Lu and Young, 2020). Moreover, the collected datasets are often limited in their generalizability to even similar agricultural domains or applications, necessitating repetition in the data collection and model development process (Ghazi et al., 2017, Thenmozhi and Reddy, 2019). This repetitive process not only adds to the overall time and monetary costs but also hinders the generalizability, efficiency, and scalability of DL frameworks in agricultural applications.
Table 1: Nomenclature
Nomenclature InstructGPT Instruction generative pretrained transformer (Ouyang et al., 2022)
ML Machine learning SAM Segment anything model (Kirillov et al., 2023)
DL Deep learning SAA+ Segment any anomaly + (Cao et al., 2023b)
AI Artificial intelligence SEEM Segment everything everywhere model (Zou et al., 2023)
GAN Generative adversarial network DMs Diffusion models
VQA Visual question answering CLIP Contrastive language-image pre-training (Radford et al., 2021)
FMs Foundation models ViT Vision transformer
[…]
VFMs Vision foundation models GLIDE Guided language to image diffusion for generation and editing (Nichol et al., 2021)
MFMs Multimodal foundation models GigaGAN Large-scale GAN for text-to-image synthesis (Kang et al., 2023)
RLFMs Reinforcement learning foundation models BLIP Bootstrapping language-image pre-training (Li et al., 2022)
BERT Bidirectional encoder representations from transformers (Devlin et al., 2018) VLP Vision-language pre-training
PaLM Pathways language model (Wei et al., 2022) KOSMOS-1 A multimodal large language model (Huang et al., 2023)
LLAMA Large language model meta AI (Touvron et al., 2023) MLLM Multimodal large language model
GLaM Generalist language Model (Du et al., 2022) DRL Deep reinforcement learning
GPT-2 Generative pretrained transformer 2 (Radford et al., 2019) DQN Deep Q-network (Mnih et al., 2015)
GPT-3 Generative pretrained transformer 3 (Brown et al., 2020) PPO Proximal policy optimization (Schulman et al., 2017)
GPT-4 Generative pretrained transformer 4 (OpenAI, 2023) Gato A aeneralist agent (Reed et al., 2022)
Chat-GPT Chat generative pretrained transformer (OpenAI, 2022) AdA Adaptive agent (Team et al., 2023a)
@@
To mitigate the aforementioned issues, many approaches have been proposed, including transfer learning (Weiss et al., 2016, Zhuang et al., 2020), few-shot learning (Wang et al., 2020, Yang et al., 2022a), label-efficient learning (Zhou, 2018, Li et al., 2023a), among others. Specifically, transfer learning has emerged as a widely adopted approach in agricultural applications and beyond (Espejo-Garcia et al., 2020, Chen et al., 2022, Abbas et al., 2021, Dang et al., 2022), which leverages DL models pre-trained on large-scale image datasets, such as ImageNet (Deng et al., 2009), Microsoft COCO (Lin et al., 2014), and PlantCLEF2022 (Goëau et al., 2022). The pre-trained models are then fine-tuned for specific tasks. For instance, 35 state-of-the-art DL models are evaluated for weed classification using transfer learning approach (Chen et al., 2022). Few-shot learning, on the other hand, focuses on training models to quickly adopt new tasks or domains using only a limited number of labeled examples, by leveraging prior knowledge or meta-learning approaches. In contrast, label-efficient learning (Li et al., 2023a) employs weak supervision or no supervision techniques to mitigate the effects of laborious and time-consuming labeling challenges. However, these methods are often pre-trained on data from a single domain and modality, which restricts their adaptability to other domains and applications (Raffel et al., 2020). The above observations highlight the importance of incorporating diverse data sources to enhance their generalization capabilities across different domains and applications.
@@
Large pre-trained models are also referred to as FMs (Bommasani et al., 2021, Moor et al., 2023, Mai et al., 2023) designed to produce a wide and general variety of outputs, undergo extensive training on vast and diverse datasets. FMs are enabled to tackle numerous downstream tasks, such as speech recognition and natural language understanding (Radford et al., 2019, Brown et al., 2020, OpenAI, 2022, Ouyang et al., 2022, Touvron et al., 2023), computer vision (Rombach et al., 2022, Ramesh et al., 2022, Kirillov et al., 2023, Kang et al., 2023), and decision making (Reed et al., 2022, Team et al., 2023a). FMs are generally trained on large-scale datasets from diverse domains and modalities through self-supervised learning (Liu et al., 2021), and are capable of handling a wide range of applications from various domains with minimum fine-tuning and no/little task-specific labeled data (Bommasani et al., 2021, Moor et al., 2023, Mai et al., 2023). For example, the Generative Pre-trained Transformer 3 language model (GPT-3, (Brown et al., 2020)) demonstrates a breakthrough in in-context learning, enabling the model to perform novel tasks without explicit training, by leveraging text explanations (known as “prompts”) containing just a few examples. Subsequently, ChatGPT (OpenAI, 2022), a derivative of GPT-3, is developed and publicized by OpenAI for chat-based interactions and has brought about a transformative evolution in natural language processing, enabling unprecedentedly engaging conversational experiences. In the computer vision domain, FMs such as the segment anything model (SAM, (Kirillov et al., 2023)), trained on over 1 billion masks on 11M licensed and privacy-respecting images, can solve a range of downstream segmentation problems with new image distributions and tasks with zero-shot generalization where a model is trained to recognize classes or categories that it has never seen during training (Yang et al., 2022a). DeepMind Ada (Team et al., 2023a), equipped with a customized Transformer architecture (Vaswani et al., 2017) using distillation with a teacher-student approach (Schmitt et al., 2018), brings FMs to reinforcement learning (RL) that generalizes to new tasks in a few-shot setting. Despite the aforementioned advances, FMs’ adoption in agriculture has received scarce attention.
As such, this study aims to explore the potential of developing and applying FMs for agricultural applications. The main contributions and the technical advancements of this paper are summarized as follows.
  • Recent FMs (Section 2) in the general computer science (CS) domain are systematically reviewed and categorized into four key categories: language FMs, vision FMs, multimodal FMs, and reinforcement learning FMs.
  • The process of developing AFMs (Figure 1) is outlined and their potential applications (Section 3) in smart agriculture are discussed, shedding light on their diverse capabilities.
  • This study also identifies and discusses the unique challenges and potential pitfalls associated with developing AFMs (Section 4), including model training, validation, and deployment.
This comprehensive examination of AFMs can serve as a valuable resource for newcomers to the field, as well as experienced researchers seeking to innovate within the agricultural space.
@@ 2 Foundation Models (FMs)
In this section, both the strengths and weaknesses of FMs are first covered, giving readers a balanced view of the topic. Then, the different kinds of FMs, their distinct characteristics, and their applicability to various domains or tasks are explored.
@@ 2.1 Pros and Cons of FMs
FMs represent a significant paradigm shift in the field of AI. In contrast to traditional AI models that are trained on domain-specific data for specialized tasks, FMs are pre-trained on a vast scale of data, spanning multiple domains and modalities (Bommasani et al., 2021, Moor et al., 2023, Mai et al., 2023). The training process allows these models to learn a broad base of knowledge and skills, much like the foundation of a building that supports all the structures above it. Once trained, these FMs can accomplish versatile tasks with just a little fine-tuning. This is analogous to the way humans transfer knowledge and skills from one area to another (Moor et al., 2023). They have led to breakthroughs in a variety of fields, from natural language processing to computer vision and beyond. The pros of FMs are summarized as follows:
@@
Given the distinct advantages of FMs over traditional AI models, they exhibit substantial promise for agricultural applications. Some compelling examples of applying LLMs and FMs in the agricultural domain can be found in Section 3. However, despite their progress, FMs are not without their unique set of challenges:
@@
For an in-depth discussion on deploying FMs in agricultural contexts, please refer to Section 4.
@@ 2.2 Categories of FMs
In this subsection, a comprehensive review of LLMs & FMs within the computer science (CS) domain is presented, and the models are categorized into four distinct types based on the primary data modality and learning paradigm they address: language FMs (LFMs), vision FMs (VFMs), multimodal FMs (MFMs), and reinforcement learning FMs (RLFMs) (see Figure 2). Readers who are familiar with these concepts or more interested in agricultural applications may skip this section and refer to Figure 2 for a concise summary.
Figure 2: Timeline of FMs.
@@ 2.2.1 Language foundation models (LFMs)
Over the past few years, the field of computational natural language has seen a complete transformation (Devlin et al., 2018, Brown et al., 2020) due to the rise of LLMs. Essentially, language modeling (Nadkarni et al., 2011) involves predicting the subsequent element in a given sequence of tokens, relying on a self-supervised objective that requires no manual labeling beyond a raw text corpus.
@@
The Bidirectional Encoder Representations from Transformers (BERT, (Devlin et al., 2018)) is one of the most prominent models, with two versions: $\text{BERT}_{\text{BASE}}$ and $\text{BERT}_{\text{LARGE}}$, containing 110M and 340M parameters respectively. By simply incorporating an additional output layer for fine-tuning, the pre-trained BERT model can be adapted to generate cutting-edge models for a variety of tasks, such as question answering and language inference. Remarkably, this can be achieved without requiring substantial modifications to the task-specific architecture, even surpassing the performance of many dedicated task-specific architectures. Subsequently, Google proposes the Pathways Language Model (PaLM, (Wei et al., 2022)) and Large Language Model Meta AI (LLAMA, (Touvron et al., 2023)), further advancing the field of natural language processing. PaLM, with 540B parameters, is designed with a dense decoder-only Transformer model trained with the Pathways system (Barham et al., 2022) to enable efficient training. PaLM surpasses the few-shot performance of prior large models, such as GLaM (Du et al., 2022), GPT-3 (Brown et al., 2020), and LLAMA (Touvron et al., 2023), on numerous very difficult tasks, such as nature language inference (Bowman et al., 2015), in-context reading comprehension (Jenkins et al., 2003), and question answering (Kwiatkowski et al., 2019).
@@
Participating in the development of large language models, OpenAI has proposed noteworthy contributions such as Generative Pretrained Transformer 2 (GPT-2, (Radford et al., 2019)), 3 (GPT-3, (Brown et al., 2020)), and 4 (GPT-4, (OpenAI, 2023)), having grown progressively more powerful, with 1.5 billion, 175 billion, and a staggering number of parameters, respectively, and setting remarkable benchmarks in language processing. ChatGPT (OpenAI, 2022), a specialized version of these models, has been extensively fine-tuned for human-like conversation. By adjusting the final output layer, the pre-trained GPT models can be tailored to deliver exceptional performance across a wide range of tasks, such as text generation, translation, and dialogue systems, outperforming many task-specific models without requiring extensive modifications to their foundational architecture. Recently, OpenAI has shown that InstructGPT (Ouyang et al., 2022) with only 1.3B parameters are much better at following users’ instructions than GPT-3, by using an existing technique called reinforcement learning from human feedback (RLHF, (Christiano et al., 2017)). More recently, the development of Claude 3111Claude 3 website: https://www.anthropic.com/news/claude-3-family. represents the cutting-edge in AI systems, encompassing undergraduate-level expert knowledge (MMLU), graduate-level expert reasoning (GPQA), basic mathematics (GSM8K), and beyond. It demonstrates nearly human-like levels of comprehension and fluency across intricate tasks, marking the forefront of general intelligence.